Calculate the compound interest on a term deposit of $30,000 at the rate of 6% annually for 3 years when the investment is compounded: a) annually. b) semi-annually. c) quarterly. d) monthly. e) daily.
1
Expert's answer
2020-01-07T08:09:41-0500

Given

P=Principal = "\$30000"

r= rate of interest ="6 \%=0.06"

t =number of time periods elapsed

n = number of times interest applied per time period


(a) compounded annually


n = 1



"A = P (1 +\frac {r}{n})^{nt}"

"A = 30000 (1 +\frac {0.06}{1})^{1\times 3} = 30000 \times (1.06)^3"

"A=30000 \times 1.191 =35,730"

"Interest = A - P =35730-30000 =\$5730"

(b) compounded semi-annually

n = 1


"A = P (1 +\frac {r}{n})^{nt}"

"A = 30000 (1 +\frac {0.06}{2})^{2\times 3} = 30000 \times (1.03)^6"

"A=30000 \times 1.194 =35,820"

"Interest = A - P =35,820-30000 =\$5820"

(c) compounded quarterly


n= 4


"A = 30000 (1 +\frac {0.06}{4})^{4\times 3} = 30000 \times (1.015)^{12}"

"A=30000 \times 1.1956 =35868"

"Interest = A - P =35,868-30000 =\$5868"



(d) compounded monthly

n=12


"A = 30000 (1 +\frac {0.06}{12})^{12\times 3} = 30000 \times (1.005)^{36}=35,900"

"Interest = A - P =35,900-30000 =\$5900"

(e) compounded daily


n=365


"A = 30000 (1 +\frac {0.06}{365})^{365\times 3} = 30000 \times (1.000164385)^{1095}"



"A =30000\times 1.1972=\$35916"


"Interest = A - P =35,916-30000 =\$5916"